# 1]FAIR at Meta

### v1

@@ 1 Introduction
We show that this behavior is due to an interaction between weight decay and learning-rate schedules that occurs only for weight layers that are affected by normalization layers such as LayerNorm Ba et al., (2016) or BatchNorm (Ioffe and Szegedy,, 2015). We present the following results:
@@ 2 Related Work
Kosson et al., (2024) view weight decay through the lens of maintaining a rotational equilibrium. D’Angelo et al., (2024) argue that weight decay’s primary effect is as a loss stabilization mechanism during multi-pass training runs. For single pass LLM training, they argue that weight decay results in larger effective learning rates and prevents blowup when training in bfloat16.
@@ 3.2 Weight decay balances the optimal learning rates of all normalized layers
Balancing the rate-of-learning across layers has been explored in prior works. Defazio and Bottou, (2022) show that for networks with ReLU activations, balancing the gradient-to-weight ratios results in the diagonal blocks of the Hessian having the same average singular value. The Fromage optimizer (Bernstein et al.,, 2020) explicitly rescales the step for each layer to control the gradient-to-weight ratio.
@@ 4.1 AdamW, not Adam, controls norm ratios
Since the two $x$ norm terms use different weighted norms, the same derivation no longer applies, and the steady-state dynamics for different layers will no longer converge to the same uniform ratio. We conjecture that this lack of balance is a major contributing factor to the performance difference between Adam and AdamW.

### v2

@@ 1 Introduction
We show that this behavior is due to an interaction between weight decay and learning-rate schedules that occurs only for weight layers that are affected by normalization layers such as LayerNorm (Ba et al.,, 2016) or BatchNorm (Ioffe and Szegedy,, 2015). We present the following results:
@@ 2 Related Work
Kosson et al., (2024) view weight decay through the lens of maintaining a rotational equilibrium. Through this framework, a link between the dynamics of layer/neuron learning rates and weight decay is established. The conclusions we draw in our work use a different analysis technique, but we come to many of the same conclusions at a high level regarding the effect of weight decay on learning. They also introduce a constrained dynamics optimization wrapper, which explicitly constrains the behavior of normalized layers, a complementary approach to the small modification to weight decay that we suggest in this work.
D’Angelo et al., (2024) argue that weight decay’s primary effect is as a loss stabilization mechanism during multi-pass training runs. For single pass LLM training, they argue that weight decay results in larger effective learning rates and prevents blowup when training in bfloat16.
@@ 3.2 Weight decay balances the optimal learning rates of all normalized layers
Balancing the rate-of-learning across layers has been explored in prior works. Defazio and Bottou, (2022) show that for networks with ReLU activations, balancing the gradient-to-weight ratios results in the diagonal blocks of the Hessian having the same average singular value. The Fromage optimizer (Bernstein et al.,, 2020) explicitly rescales the step for each layer to control the gradient-to-weight ratio. Kosson et al., (2024) consider effective learning rates as the size of the angular update for normalized layers. Through this framework they also show that a balancing effect is induced by weight-decay.
@@ 4.1 AdamW, not Adam, controls norm ratios
Since the two $x$ norm terms use different weighted norms, the same derivation no longer applies, and the steady-state dynamics for different layers will no longer converge to the same uniform ratio. We conjecture that this lack of balance is a major contributing factor to the performance difference between Adam and AdamW. This conclusion complements the result of Kosson et al., (2024), who also consider the differences between Adam and AdamW as arising from unbalanced layer-wise effective learning rates. They consider the size of angular updates from the two methods, and show that Adam results in unbalanced angular updates when the gradient norm varies between layers.